The force of interest delta(t) is given by ln(1.5) for all non-negative t.

Find the accumulation function at 4, a(4).

A.

5.0625

B.

6.7420

C.

5.2432

D.

The correct answer does not appear here

E.

4.0625
Answer #1

The realtionship between force of interest \delta (k) and accumulation function a(t) is given by

a(t) = es(s)ds

Let us assume that \delta (t) is a constant function say, \delta (t) = c, for all t where c is the constant.

a(t) = ecds

a(t) = ecds

Then, a(t) = ect since (7 ds Jo = s[t] - s[0] = t - 0 = t)

Solving we have \delta (t) = ln(1.5) for all non negative 't'

This implies that a(t) = e[ln(1.5)]^t

a(t) = 1.5t

Solving for the given case where t = 4, we get a(t) = 5.0625

Option (A)

Cheers !!!